# Sonnet: Spectral Operator Neural Network for Multivariable Time Series Forecasting

### v1

@@ C.1 Baseline MTS forecasting models
C.2 Attention mechanisms used for time series forecasting
Variable Deformable Attention Block (VDAB). Introduced in DeformTime [43], VDAB adaptively attends to relevant information across time and variables by applying cross attention over the original and deformed input. This enables to capture complex inter-variable dependencies in MTS forecasting.
Frequency Enhanced Decomposition (FED). Employed in FEDformer [62], FED decomposes time series into trend and seasonal components, applying attention mechanisms to each, thereby capturing both periodic patterns and time-delay dependencies effectively. We note that we do not compare to the complete forecasting model of FEDformer as it has been outperformed by most of our selected baseline models. Crossformer, iTransformer, PatchTST, ModernTCN, DLinear, and Samformer have all provided direct comparison with FEDformer in their original papers. Meanwhile, the ones that do not provide a direct comparison with FEDformer, namely TimeXer and DeformTime, have both outperformed PatchTST and iTransformer on MTS forecasting tasks by a large margin. Given Sonnet’s competitive performance over all the aforementioned (baseline) models, we expect it to yield superior performance compared to FEDformer.
FNet. The attention proposed in FNet [21] replaces the commonly used self-attention mechanism in a transformer with a parameter-free Fourier transform. Specifically, it maps the input sequence to the frequency domain and retains only the real part of the result. This approach enables effective mixing between input tokens along the time dimension without adding learnable parameters. FNet is not included as a baseline model in the main results as it is not a model designed for forecasting purposes. Additionally, we show in sections 4.3 and 4.4 that it gives a similar performance to the vanilla transformer, which has been outperformed.
Naïve Attention. The commonly used attention module was first proposed by Vaswani et al. [48] as a core component of the transformer model. The attention weights are computed using scaled dot product operations between query and key embeddings, facilitating the model’s ability to focus on relevant parts of the input sequence.
Table S2: ILI rate forecasting results in US Region 2 (ILI-US2) with web search frequency time series as exogenous variables. We assess performance on $4$ test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
@@ C.2 Attention mechanisms used for time series forecasting
Appendix D Supplementary experiment settings
D.1 Hyperparameter settings
The look-back window $L$ is set differently depending on the forecasting task. For the ETTh1, ETTh2 data sets, $L$ is set to $336$ time steps for all $4$ different forecasting horizons $H=\{96,192,336,720\}$, based on prior work [34]. In ELEC experiments, we set $L=168$ hours (a week, given the strong weekly periodicity in these time series) for all $H=\{12,24,36\}$. For ENER, $L$ is set equal to the forecasting horizon, i.e. $L=H=\{24,48,72,168\}$. For the WEA tasks, we use a look-back window $L=\{28,28,56,240\}$ time steps for $H=\{4,12,28,120\}$, correspondingly. This means that we use the past $7$ days to conduct short-term forecasts ($1$ and $3$ days ahead), and the past $2H$ days for longer-term forecasts. For the ILI tasks, we set $L=\{28,28,56,56\}$ days respectively for forecasting horizons $H=\{7,14,21,28\}$ days ahead following prior work [43].
Table S3: ILI rate forecasting results in US Region 9 (ILI-US9) with web search frequency time series as exogenous variables. We assess performance on $4$ test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
@@ D.1 Hyperparameter settings
D.2 Hyperparameter settings specific to Sonnet
[…]
D.4 Hyperparameter validation
[…]
Table S4: WEA forecasting results (temperature at $850$ hPa) in London (WEA-LD). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
@@ D.4 Hyperparameter validation
D.5 Data normalisation
We apply z-score normalisation to the data of all forecasting tasks. This standardises endogenous and exogenous variables in a training set to have a zero mean and unit standard deviation. Normalisation is transferred from the training data to the test data. Data is reverted back to the original scale when computing prediction errors. Additionally, for the ETT, ENER, and ELEC data sets, we adopt variable normalisation [19] over the target variable within the input’s look-back window following prior work [34, 43].
Table S5: WEA forecasting results (temperature at $850$ hPa) in New York (WEA-NY). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
@@ D.5 Data normalisation
D.6 sMAPE definition for the weather (temperature) prediction tasks
We note that for the WEA tasks, the unit of the temperature time series sequence is originally captured in Kelvin.99900 degrees in Celsius is equal to $273.15$ Kelvin. Directly computing the sMAPE using Kelvin units will lead to a relatively small error (MAEs are not affected, but sMAPEs will be significantly lower and harder to display in an accessible way). To address this, we subtract the minimum value of the target variable across the test set, and then add a constant number $a=30$ to both the ground truth and the predicted value to avoid division by 00. This alters the formula of sMAPE as follows:
$\text{sMAPE}(\mathbf{y},\hat{\mathbf{y}})=\frac{100}{L}\sum_{t=1}^{L}\frac{|y_{t}-\hat{y}_{t}|}{\left(|y_{t}-\min(\mathbf{y})+a|+|\hat{y}_{t}-\min(\mathbf{y})+a|\right)/2}\times 100\%\,\,.$ (S1)
D.7 Random seed initialisation
We use a fixed seed equal to ‘42’ during training for all forecasting tasks, except for the ETTh1 and ETTh2 data sets. For these data sets, we follow prior work [34, 58] and use ‘2021’, with the exception of iTransformer, which uses ‘2023’ in line with its official configuration.
Table S6: WEA forecasting results (temperature at $850$ hPa) in Hong Kong (WEA-HK). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
@@ D.7 Random seed initialisation
D.8 Replacing attention modules with MVCA
Table S7: WEA forecasting results (temperature at $850$ hPa) in Singapore (WEA-SG). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.
@@ D.8 Replacing attention modules with MVCA
D.9 Additional information for Sonnet’s ablation study
Information in this section supplements section 4.5 in the main paper.
$\neg$ MLP: This denotes that we do not add the residual connection with the two-layer perceptron to obtain the attention output. Instead of computing $\mathbf{O}_{m}=\mathbf{O}_{r}+\texttt{MLP}\left(\mathbf{O}_{r}\right)$, we directly have $\mathbf{O}_{m}=\mathbf{O}_{r}$. In this case, the inter-variable dependency via the MLP is omitted. Although some inter-variable interactions are partially (no learnable parameter is included in this process) captured when transforming the input into the frequency space along the feature dimension.
Table S8: WEA forecasting results (temperature at $850$ hPa) in Cape Town (WEA-CT). We assess performance on $3$ test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded and second best are underlined.

### v2

@@ H Model 2016 2017 2018 2019 Average
Table S1: ILI rate forecasting results in England (ILI-ENG) with web search frequency time series as exogenous variables. We assess performance on 44 test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
A.4 Weather forecasting on selected locations
We form the weather forecasting (WEA) data set using climate information extracted from the WeatherBench repository (rasp2020weatherbench), covering the period from 1979 to 2018.666WeatherBench repository,
dataserv.ub.tum.de/index.php/s/m1524895
Data is sampled from a uniform spatial grid over the globe with a resolution of $5.625^{\circ}$ in both latitude and longitude. Each grid point represents a distinct geographical location. To capture different climatic conditions, we select five spatially diverse cities from the grid: London (UK), New York (US), Hong Kong (China), Cape Town (South Africa), and Singapore. Among these, London, New York, and Hong Kong are located in the northern hemisphere, with Hong Kong being in the subtropical zone; Singapore lies near the equator, and Cape Town is in the southern hemisphere. The two cities that are closer to the equator (Hong Kong and Singapore) do not have strong seasonal temperature fluctuations over the year. Consequently, we expect models to yield lower prediction errors at these two locations. For each city, climate data is sampled from the nearest grid point to its coordinates. [^7] We obtain the longitude and latitude of each city using OpenStreetMap (wiki.openstreetmap.org). Specifically, we obtain the absolute distance in both latitude and longitude between each city and all grid points, and select the grid point with the smallest distance along both axes. The forecasting target is the $850$ hPa temperature in Kelvin (T850) at the central grid point, [^8] Further information about T850 is provided at
charts.ecmwf.int/products/medium-z500-t850.
which is an important climate indicator that has been widely used in climate research (scherrer2004analysis; hamill2007ensemble). To form exogenous inputs, we include four other climate variables at each selected grid point following prior work (nguyen2023climax; verma2024climode). These are the $500$ hPa geopotential (Z500), $2$-meter air temperature (t2m), $10$-meter zonal wind (u10), and $10$-meter meridional wind (v10). We also include these $4$ variables and T850 from the surrounding eight grid points, forming a $3\times 3$ spatial window centred on each selected grid, resulting in $44$ exogenous variables in total. The data is resampled to maintain a temporal resolution of $6$ hours similarly to verma2024climode.
We consider $3$ test seasons: years 2016, 2017, and 2018 (Jan. 1 to Dec. 31). Each test year contains 1,4641{,}464, 1,4601{,}460, and 1,4601{,}460 time steps, respectively. The year immediately preceding a test year is used for validation (2015, 2016, and 2017). Validation year 2015 contains 1,4601{,}460 time steps. The training sets include data from January 1, 1980, up to the date preceding the corresponding validation year, resulting in 51,13651{,}136, 52,59652{,}596, and 54,06054{,}060 time steps for the $3$ test seasons.
@@
Table S2: ILI rate forecasting results in US Region 2 (ILI-US2) with web search frequency time series as exogenous variables. We assess performance on 44 test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
A.5 Influenza-like-illness rate prediction
We construct an ILI forecasting data set for three regions: England (UK), and US HHS Regions 2 and 9, following the setup in prior work (shu2025deformtime). Ground truth ILI rates are obtained from the Royal College of General Practitioners (RCGP) in England and the Centers for Disease Control and Prevention (CDC) for the US regions. Weekly ILI rates are linearly interpolated to a daily resolution. Specifically, we assume the ILI rates are reported corresponding to the middle day of each week, which is Wednesday for the US regions and Thursday for England (shu2025deformtime). We also incorporate the ILI rate reporting delay of syndromic surveillance systems; in England there is a lag of $\delta=7$ days and $\delta=14$ days for the US regions (wagner2018added; reich2019accuracy).
We use web search activity data (daily frequency time series of different search terms), obtained from the Google Health Trends API, as the exogenous indicators. We follow the setup of DeformTime (shu2025deformtime) to form the set of search queries for our experiments. [^9] Search queries can be obtained from github.com/ClaudiaShu/DeformTime. Note that for the US regions, search frequencies are collected at the state level, and the query data for each region is obtained using weighted averages based on the population of the states inside each region (New Jersey/New York is $0.32/0.68$ for Region 2 and Arizona/California is $0.16/0.84$ for Region 9).
We evaluate forecasting accuracy using $4$ test seasons (2015/16, 2016/17, 2017/18, and 2018/19) in England and the US regions, each corresponding to a 12-month period that encompasses at least one seasonal flu outbreak. For England, each season begins on September $1$ and ends on August $31$ of the next year, while for the US regions, it is from August $1$ to July $31$ of the following year. For each test season, models are trained on the preceding $9$ seasons. Validation sets are extracted from the last three seasons in the training data. Each validation set covers $180$ days, containing three $60$-day segments that capture the onset (from the last season), peak (from the penultimate season), and outset (the 33rd to last season) phases of an influenza season. The onset is defined as the first point followed by two weeks above a predefined ILI threshold; the peak is the point of the highest ILI rate; and the outset is the last point followed by two weeks above the threshold. These selected points are then considered as the $30$th day in a $60$-day validation period.
@@
Table S3: ILI rate forecasting results in US Region 9 (ILI-US9) with web search frequency time series as exogenous variables. We assess performance on 44 test seasons (2015/16 to 2018/19) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
Appendix B Further information about MVCA
In Equation 2, the cross-spectral and power-spectral density matrices are obtained using the frequency domain embeddings $\mathbf{Q}_{f}$, $\mathbf{K}_{f}\in\mathbb{C}^{L\times\ell}$. The obtained matrices $\mathbf{P}_{qq}=\mathbf{Q}_{f}\odot\mathbf{Q}_{f}^{\ast}$ and $\mathbf{P}_{kk}=\mathbf{K}_{f}\odot\mathbf{K}_{f}^{\ast}$ are both containing real numbers because multiplying a complex number (e.g. $z=a+ib$) by its conjugate (aiba-ib) results in its squared magnitude ($a^{2}+b^{2}$), which is a real number.
Appendix C Baseline forecasting models and attention mechanisms
C.1 Baseline MTS forecasting models
DeformTime (shu2025deformtime) is specifically designed for multivariable forecasting. It uses deformable attention to capture the dependencies across time and covariates, reaching competitive performance on MTS forecasting tasks. Based on our evaluation, DeformTime is the best-performing baseline.
ModernTCN (luo2024moderntcn) applies convolution along the temporal and variable axes. By leveraging large receptive fields, it effectively captures both intra-variable and inter-variable dependencies.
TimeXer (wang2024timexer) is a transformer-based model that focuses on multivariable forecasting. It first embeds the exogenous variables along the sequence dimension, and then captures the exogenous information through a cross-attention mechanism.
Samformer (ilbert2024samformer) first embeds the input along the sequence dimension and then applies naïve attention along the feature dimension. It conducts forecasts using a linear layer. We also include Samformer as one of the base models that deploy naïve attention.
iTransformer (liu2024itransformer) embeds the input along the sequence dimension and deploys multiple attention layers along the variable dimension. We also include iTransformer as one of the base models.
@@ H Model 2016 2017 2018 Average
Table S4: WEA forecasting results (temperature at $850$ hPa) in London (WEA-LD). We assess performance on 33 test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
PatchTST (nie2023time) segments time series data into patches of fixed size, treating them as input tokens to a transformer. It processes each variable independently, enabling efficient modelling of long-term dependencies. PatchTST is also used as one of the base models.
Crossformer (zhang2023crossformer) captures both inter- and intra-variable dependencies using a two-stage attention mechanism. Each stage is dedicated to one dependency type. The information is captured using time series patches.
DLinear (zeng2023are) uses a simpler neural network architecture (feedforward layers). It decomposes each input series into trend and seasonal components, which are then forecasted independently using linear layers. Each variable is modelled separately, using its own historical values for prediction. Based on our evaluation, DLinear is the worst-performing model in MTS forecasting. It does not outperform persistence in most experiments (Table 1). This corroborates findings in related work (shu2025deformtime).
Naïve forecasting models. We include simple baseline models as reference points for performance comparison. We use either seasonal persistence or standard persistence, depending on whether a target variable (and task) is characterised by strong seasonality or not. A standard persistence model (naïve baseline) uses the last observed value of the target variable as the prediction (forecasting outcome). The seasonal persistence model uses the value from the same point in the previous seasonal cycle. In our case, the electricity consumption data (ELEC) has a strong seasonal trend, with seasonal persistence outperforming the standard one by a large margin (results not shown). Therefore, we use the seasonal persistence model for the ELEC task, while for the other tasks, we retain standard persistence.
@@
Table S5: WEA forecasting results (temperature at $850$ hPa) in New York (WEA-NY). We assess performance on 33 test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
C.2 Additional information on adapting baseline models for Multivariable forecasting tasks
Amongst the selected baselines, DeformTime (shu2025deformtime) and TimeXer (wang2024timexer) are specifically designed for multivariable forecasting tasks. For the rest of the models that were designed for multivariate forecasting, we follow the adaptation strategy from DeformTime and TimeXer, which conduct multivariate forecasting and evaluate the model performance on the target variable. Given that DeformTime and TimeXer have demonstrated their effectiveness across a wide range of forecasting benchmarks, we believe this evaluation provides a fair and consistent comparison against SOTA multivariate forecasting models.
C.3 Attention mechanisms used for time series forecasting
Variable Deformable Attention Block (VDAB). Introduced in DeformTime (shu2025deformtime), VDAB adaptively attends to relevant information across time and variables by applying cross attention over the original and deformed input. This enables to capture complex inter-variable dependencies in MTS forecasting.
Frequency Enhanced Decomposition (FED). Employed in FEDformer (zhou2022fedformer), FED decomposes time series into trend and seasonal components, applying attention mechanisms to each, thereby capturing both periodic patterns and time-delay dependencies effectively. We note that we do not compare to the complete forecasting model of FEDformer as it has been outperformed by most of our selected baseline models. Crossformer, iTransformer, PatchTST, ModernTCN, DLinear, and Samformer have all provided direct comparison with FEDformer in their original papers. Meanwhile, the ones that do not provide a direct comparison with FEDformer, namely TimeXer and DeformTime, have both outperformed PatchTST and iTransformer on MTS forecasting tasks by a large margin. Given Sonnet’s competitive performance over all the aforementioned (baseline) models, we expect it to yield superior performance compared to FEDformer.
FNet. The attention proposed in FNet (lee-thorp2022fnet) replaces the commonly used self-attention mechanism in a transformer with a parameter-free Fourier transform. Specifically, it maps the input sequence to the frequency domain and retains only the real part of the result. This approach enables effective mixing between input tokens along the time dimension without adding learnable parameters. FNet is not included as a baseline model in the main results as it is not a model designed for forecasting purposes. Additionally, we show in Sections 4.3 and 4.4 that it gives a similar performance to the vanilla transformer, which has been outperformed.
Naïve Attention. The commonly used attention module was first proposed by vaswani2017attention as a core component of the transformer model. The attention weights are computed using scaled dot product operations between query and key embeddings, facilitating the model’s ability to focus on relevant parts of the input sequence.
@@
Table S6: WEA forecasting results (temperature at $850$ hPa) in Hong Kong (WEA-HK). We assess performance on 33 test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
Appendix D Supplementary experiment settings
D.1 Hyperparameter settings
The look-back window $L$ is set differently depending on the forecasting task. For the ETTh1, ETTh2 data sets, $L$ is set to $336$ time steps for all $4$ different forecasting horizons $H=\{96,192,336,720\}$, based on prior work (nie2023time). In ELEC experiments, we set $L=168$ hours (a week, given the strong weekly periodicity in these time series) for all $H=\{12,24,36\}$. For ENER, $L$ is set equal to the forecasting horizon, i.e. $L=H=\{24,48,72,168\}$. For the WEA tasks, we use a look-back window $L=\{28,28,56,240\}$ time steps for $H=\{4,12,28,120\}$, correspondingly. This means that we use the past $7$ days to conduct short-term forecasts ($1$ and $3$ days ahead), and the past $2H$ days for longer-term forecasts. For the ILI tasks, we set $L=\{28,28,56,56\}$ days respectively for forecasting horizons $H=\{7,14,21,28\}$ days ahead following prior work (shu2025deformtime).
@@
Table S7: WEA forecasting results (temperature at $850$ hPa) in Singapore (WEA-SG). We assess performance on 33 test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
D.2 Hyperparameter settings specific to Sonnet
For Sonnet, the parameter $\alpha$ that controls the dimensionality of the embeddings of the endo- and exogenous variables is chosen from $\alpha\in\{0,.1,.25,.75\}$. The number of atoms, $K$, is selected from the set $K\in\{8,16,32\}$. The hidden dimension $d$ of Sonnet is set to be $64$.
In the ILI rate forecasting task, the number of exogenous variables is decided using a learnable threshold $\tau$ based on linear correlation (with the endogenous/target variable in the training data), following DeformTime (shu2025deformtime). We select $\tau$ from $\{.3,.4,.5\}$ for the US regions, and from $\{.05,.1,.2,.3,.4,.5\}$ for England. An increase in the correlation threshold leads to a reduction in the number of selected queries. Consequently, the number of search queries selected by the models ranges from $13$ to $51$ for US regions, and from $80$ to $752$ for England. For the models that require an excessive amount of GPU memory for larger sets of exogenous variables, i.e. ModernTCN, PatchTST and Crossformer, we restricted $\tau\in\{.3,.4,.5\}$ for all regions.
D.4 Hyperparameter validation
For the ETT forecasting tasks, the batch size and the hidden dimension $d$ are both selected from $\{16,32,64\}$. For all tasks, Adam’s initial learning rate is selected from $\{2,1,.5,.2,.1,.05\}\times 10^{-3}$. The learning rate is decayed to 0 with a linear decay scheduler applied after every epoch. If a model structure contains a dropout parameter, its value is selected from $\{0,.1,.2\}$. Hyperparameter values are determined using grid search. Training stops early when the validation error of the target variable over time $\{t+1,\cdots,t+H\}$ does not decrease further (compared to its lowest value) for $5$ consecutive epochs. We then use the model with the lowest validation loss.
@@
Table S8: WEA forecasting results (temperature at $850$ hPa) in Cape Town (WEA-CT). We assess performance on 33 test periods (years 2016, ’17, ’18) using linear correlation ($r$), MAE, and sMAPE ($\epsilon\%$). Best results are bolded, second best underlined.
D.5 Data normalisation
We apply z-score normalisation to the data of all forecasting tasks. This standardises endogenous and exogenous variables in a training set to have a zero mean and unit standard deviation. Normalisation is transferred from the training data to the test data. Data is reverted back to the original scale when computing prediction errors. Additionally, for the ETT, ENER, and ELEC data sets, we adopt variable normalisation (kim2022reversible) over the target variable within the input’s look-back window following prior work (nie2023time; shu2025deformtime).
Model Variant $H=7$ $H=14$ $H=21$ $H=28$
ENG US2 US9 ENG US2 US9 ENG US2 US9 ENG US2 US9
iTransformer
26.38 23.24 18.57 36.67 28.17 22.44 48.93 30.03 24.11 55.35 36.75 31.05
FNet (lee-thorp2022fnet) 30.16 26.99 22.27 41.28 31.12 26.29 57.16 36.48 29.69 64.26 41.06 35.00
FED (zhou2022fedformer) 40.21 32.58 26.25 48.37 37.86 30.68 67.76 50.94 40.97 73.86 54.82 44.33
VDAB (shu2025deformtime) 27.46 24.55 19.21 38.10 30.39 23.29 48.19 32.37 23.07 54.16 35.39 26.97
SamFormer
28.31 24.21 19.39 36.63 30.16 24.50 54.41 31.42 24.22 60.78 36.8 27.95
FNet (lee-thorp2022fnet) 30.01 27.94 21.95 40.33 33.06 27.33 53.16 36.21 31.30 61.64 42.28 34.68
FED (zhou2022fedformer) 40.18 32.35 26.34 48.91 37.65 30.84 67.79 49.99 41.26 73.17 53.91 44.85
VDAB (shu2025deformtime) 25.98 21.88 17.05 36.84 27.48 20.95 46.52 30.78 22.29 57.09 33.49 25.63
PatchTST
27.61 24.51 19.34 37.76 30.11 24.09 51.11 36.70 29.40 59.60 42.61 33.35
FNet (lee-thorp2022fnet) 31.93 28.81 22.14 44.15 33.77 25.64 58.17 38.11 26.46 60.51 41.48 29.92
FED (zhou2022fedformer) 40.21 32.59 26.25 50.26 37.88 30.76 67.85 50.89 41.16 74.08 54.61 44.42
VDAB (shu2025deformtime) 26.88 22.99 17.99 38.08 30.99 20.64 50.49 32.45 23.90 60.16 34.89 25.61
Table S9: Performance (average sMAPE across the $4$ test seasons) of iTransformer, Samformer, and PatchTST on the ILI forecasting tasks (ILI-ENG/US2/US9) with different modifications to the naïve attention mechanism. $\neg$ Attn’ denotes the removal of the residual attention module, and FNet / FED / VDAB refer to using the attention modules proposed in FNet, FEDformer, and DeformTime, respectively. This table supplements Table 2. Best results are bolded, second best underlined.
D.6 sMAPE definition for the weather (temperature) prediction tasks
We note that for the WEA tasks, the unit of the temperature time series sequence is originally captured in Kelvin.1010100 degrees in Celsius is equal to $273.15$ Kelvin. Directly computing the sMAPE using Kelvin units will lead to a relatively small error (MAEs are not affected, but sMAPEs will be significantly lower and harder to display in an accessible way). To address this, we subtract the minimum value of the target variable across the test set ($\xi=\min(\mathbf{y})$), and then add a constant number $a=30$ to both the ground truth and the predicted value to avoid division by 0. This alters the formula of sMAPE as follows:
$\mathrm{sMAPE}(\mathbf{y},\hat{\mathbf{y}})=\frac{100}{L}\sum_{t=1}^{L}\frac{2|y_{t}-\hat{y}_{t}|}{\left(|y_{t}-\xi|+|\hat{y}_{t}-\xi|+2a\right)}\,.$ (S1)
D.7 Random seed initialisation
We use a fixed seed equal to ‘42’ during training for all forecasting tasks, except for the ETTh1 and ETTh2 data sets. For these data sets, we follow prior work (nie2023time; zeng2023are) and use ‘2021’, with the exception of iTransformer, which uses ‘2023’ in line with its official configuration.
$\neg$ Coher $\neg$ MLP $\neg$ MVCA $\neg$ Embed $\neg$ Koop
$H$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$ MAE $\Delta\%$
ILI-ENG
7 1.422 –3.8 1.531 +3.5 1.591 +7.5 1.633 +10.4 1.486 +0.4
14 2.110 +9.7 2.171 +12.9 2.235 +16.2 2.066 +7.5 2.079 +8.2
21 2.880 +14.7 2.614 +4.1 2.794 +11.3 2.647 +5.4 2.630 +4.8
28 3.420 +24.4 3.114 +13.3 3.274 +19.2 3.091 +12.5 3.062 +11.4
ILI-US2
7 0.414 +8.8 0.411 +8.0 0.432 +13.4 0.416 +9.4 0.420 +10.5
14 0.512 +14.0 0.507 +13.0 0.528 +17.5 0.530 +17.9 0.496 +10.5
21 0.609 +14.3 0.596 +12.0 0.671 +26.0 0.573 +7.6 0.590 +10.8
28 0.623 +7.7 0.662 +14.4 0.666 +15.0 0.606 +4.8 0.634 +9.5
ILI-US9
7 0.272 +1.9 0.302 +13.1 0.277 +4.0 0.278 +4.4 0.280 +4.8
14 0.318 +13.4 0.334 +19.0 0.331 +17.8 0.315 +12.1 0.315 +12.3
21 0.358 +12.5 0.326 +2.7 0.393 +23.6 0.311 –2.1 0.319 +0.4
28 0.428 +16.4 0.345 –6.2 0.401 +9.2 0.359 –2.4 0.348 –5.4
WEA-CT
4 1.663 +2.4 1.643 +1.2 1.636 +0.7 1.678 +3.3 1.650 +1.6
12 3.591 +1.3 3.574 +0.9 3.589 +1.3 3.572 +0.8 3.568 +0.7
28 3.783 +1.5 3.760 +0.9 3.862 +3.6 3.765 +1.0 3.754 +0.7
120 3.769 +0.9 3.764 +0.7 3.768 +0.8 3.766 +0.8 3.767 +0.8
WEA-HK
4 0.639 +0.0 0.639 +0.1 0.631 –1.2 0.655 +2.6 0.639 +0.1
12 1.290 +4.4 1.261 +2.0 1.302 +5.4 1.278 +3.4 1.273 +3.1
28 1.475 +4.3 1.467 +3.8 1.477 +4.5 1.467 +3.8 1.463 +3.5
120 1.635 +5.7 1.627 +5.2 1.620 +4.7 1.599 +3.4 1.604 +3.7
WEA-LD
4 1.754 +1.8 1.767 +2.6 1.742 +1.1 1.787 +3.7 1.755 +1.9
12 2.991 +1.1 3.007 +1.6 3.001 +1.4 3.019 +2.0 2.993 +1.1
28 3.313 +3.0 3.281 +2.0 3.328 +3.5 3.326 +3.4 3.268 +1.6
120 3.385 +4.3 3.347 +3.1 3.375 +4.0 3.359 +3.5 3.345 +3.1
WEA-NY
4 1.296 +1.9 1.294 +1.8 1.281 +0.7 1.316 +3.5 1.284 +1.0
12 2.447 –0.0 2.464 +0.7 2.468 +0.9 2.467 +0.8 2.460 +0.5
28 2.751 +2.9 2.708 +1.3 2.764 +3.4 2.739 +2.4 2.723 +1.8
120 2.856 +5.3 2.777 +2.4 2.938 +8.3 2.840 +4.7 2.801 +3.2
WEA-SG
4 0.342 –0.8 0.345 +0.1 0.343 –0.5 0.348 +1.1 0.343 –0.4
12 0.423 +1.6 0.419 +0.8 0.426 +2.4 0.423 +1.6 0.422 +1.4
28 0.489 +5.2 0.476 +2.2 0.474 +1.9 0.475 +2.2 0.476 +2.3
120 0.500 +3.5 0.494 +2.3 0.503 +4.1 0.497 +2.8 0.495 +2.5
Table S10: Ablation study results across the ILI (4 test seasons) and WEA tasks (3 test years). MAE is shown for each ablated variation ($\neg$ denotes removal of a component). $\Delta\%$ is the MAE percentage of difference compared to Sonnet’s MAE (which can be seen in Table 1). MAE values are rounded to 3 decimals for spacing. Grey background denotes a performance drop (increased MAE) compared to Sonnet.
D.8 Replacing attention modules with MVCA